When sulfuric acid reacts with sodium bicarbonate, carbon dioxide gas is released. If 2.0 L of CO2 gas is collected at standard temperature and pressure, how many grams of sodium bicarbonate we consumed? (use ideal gas law - R = 0.0821 l*atm/mol*k) H2So4 (aq) + 2NaHCO3 (aq) -> Na2SO4 (aq) + 2H2O (I) + 2CO3 (g)

a) 7.5 g

b) 8.7 g

c) 3.9 g

d) 4.7 g
Answer #1

H_2SO_4(aq)+2NaHCO_3(aq)\rightarrow Na_2SO_4(aq)+ 2H_2O(l)+2CO_2(g)

From the reaction, it is clear that for every 2 moles of sodium bicarbonate consumption, 2 moles of carbon dioxide gas is formed.

Hence the stoichiometric ratio of NaHCO_3:CO_2=1:1

1mol or 22.4 L CO_2 \equiv 1 mol of NaHCO3

2 L CO_2 \equiv\frac{2*1}{22.4}=0.089\: mol\:NaHCO_3

Now,

number\:of\:moles=\frac{given\:mass}{molar\;mass}

0.089mol=\frac{given\:mass}{84g/mol}

0.089mol*84g/mol=given\:mass

given\:mass=7.47g\approx 7.5g

option a) 7.5g of sodium bicarbonate is consumed